The coins in the store's cash register total $12.50. The cash register contains only nickels, dimes and quarters. There are twice as many dimes as nickels. There are also twoice as many quarters as dimes. How many dimes are in the cash register?
Answer by dukesmart(20)

Let put Number of Nickel = N
Number of Dime = D
Number of Quarter= Q
and we know : D= 2N ; Q= 2D
Therefore we can establish the equation as follows:
N*5 + D*10 +Q*25 = 1250 ( notice: $12.50 = 1250 cents)
D/2 *5 + D*10 + 2D*25 =1250
5D + 20D + 100D = 2500 ( result is by multiplying the equation to 2)
125D = 2500
D = 20 dimes